How can the number 3.571428 be written as a ratio of two integers??

Thanks for help !!
Answer #1

Hi..

Decimals can be thought of as the calculated form of a ratio, i.e. a fraction.

For example: .5 is just another way of saying 1:2, or 1/2

Put in mathematical terms: mathImg.php?%5Cdisplaystyle%7B0.5%7D%3D%

If we multiply both sides by the fraction's denominator, we get the numerator.

2 x (0.5) = 1

So: Denominator x Decimal form = Numerator

We need to find the Denominator and Numerator; our only clue is that these will be integers.

What we need to do is multiply the decimal by a series of integers that will progressively eliminate the lowest; basically, multiply so that the last number will become a number ending in 0. For example, the last digit in the given decimal is an 8: multiplying 8 by 5 will give us 40, which allows us to get rid of the millionths place and get that much closer to a whole number.

3.571428 x 5 = 17.85714

17.85714 x 5 = 89.2857

89.2857 x 10,000 = 892,857 (the numerator).

5x5x10,000 = 250,000 (the total number we divided by).

So 3.571428 = 892,857 / 250,000.